OpenStudy (anonymous):

i need help with this 1. 1 √16 2. 2 √121 3. 3 √81 4. 4 √144 5. 5 √4 6. 6 √169 7. 7 √49 8. 8 √9 9. 9 √100 10. 10 √36 11. 11 √64 12. 12 √25 13. 13 √1

OpenStudy (anonymous):

1. =4 2. = 11 3. =9 Do you understand why the square root of 16 equals 4

OpenStudy (anonymous):

it is because 4*4= 16. Just like when 11*11= 121 and 9*9=81